# ChemMLLM: Chemical Multimodal Large Language Model

### v1

@@ 6 Conclusion
This paper has proposed ChemMLLM, a chemical multimodal large language model that handles molecule comprehension and generation across three modalities (text, SMILES string, molecule image). By jointly modeling text, SMILES strings, and molecule images, ChemMLLM enables seamless cross-modal comprehension and generation, outperforming state-of-the-art MLLMs and specialized chemical LLMs across a range of tasks. Also, we design five cross-modal chemistry tasks and curate datasets, providing a valuable resource for multimodal AI in chemistry. The experimental results demonstrate ChemMLLM’s strong performance, highlighting its potential for real-world drug and material discovery.
Limitations
Ethics Statement
The development and application of chemical AI models, such as ChemMLLM, raise important ethical considerations. We ensure that all datasets used in this work are sourced from publicly available, non-sensitive chemical data and do not involve personal or private information. The model is designed for scientific research purposes, including drug discovery and materials science, with the aim of advancing chemical understanding and innovation. We advocate for responsible use of AI in chemistry and encourage transparency, reproducibility, and fairness in future deployments of such technologies.
@@ Appendix A Data Curation Details
(1) img2caption: The dataset used for this task is sourced from chebi-20 Edwards et al. (2022)and mol-instruct Fang et al. (2023). The original datasets contain SMILES-caption pairs. We utilize RDKit Landrum et al. (2006) to transfer the SMILES into 256×256 image to form image-caption pairs. For mol-instruct, we filter captions shorter than 150 words to screen clearer descriptions. We only use the test set of chebi-20 as the test set and the partitioning of the dataset is the same as chebi-20 Edwards et al. (2022).
(2) img2property: The dataset used for this task is sourced from PubChem Kim et al. (2021). For one compound, there are three fileds related to it’s SMILES, i.e., "PUBCHEM_SMILES" which means The Simplified Molecular Linear Input Specification (SMILES) for compounds, it is a string used to represent the chemical structure, "PUBCHEM_OPENEYE_CAN_SMILES" which means Canonical SMILES generated using the OpenEye tool Software (2023) and "PUBCHEM_OPENEYE_ISO_SMILES" which means The isomer SMILES generated using the OpenEye tool. The processing steps are as follows:
  1. 1.
    Extract all the three fields from PubChem and remove the duplicate SMILES.
  2. 2.
@@
    Use the Draw.MolToImage() function in RDKit Landrum et al. (2006) to transfer SMILES into images.
  3. 3.
    Sample 100,000 SMILES for this work.
  4. 4.
    Choose 7 important properties as the prediction objective, i.e., MW, LogP, TPSA, HBD, HBA, RB and QED. Then use RDKit to calculate the 7 properties for each sampled SMILES.
  5. 5.
    Use natural language templates to integrate properties into natural language to form the final image-property answer pairs.
  6. 6.
    Divide the dataset into training and test set by the ratio of 95:5.
@@
(3) img2smiles: The dataset and the construction steps are the same as the img2property task, the difference is that this task only apply templates for SMILES to construct image-SMILES answer pairs.
@@
(4) property2img: This task is a reverse task of img2property. By swapping the question and answer of the img2property dataset, we construct property prompt-image pairs for property2img task.
(5) img2img: The dataset for this task is sourced from TDC Huang et al. (2021). The original dataset contains SMILES-SMILES pairs. The previous molecule’s LogP is lower while the latter molecule’s LogP is higher. We also use Draw.MolToImage() function in RDKit Landrum et al. (2006) to transfer them into image-image pairs. For data partitioning, we divide the training set and test set by the ratio of 9:1.
All the data used in this paper is publicly available.
@@ Appendix B Mathematical Notations
For ease of understanding, we list key mathematical notations in Table 10 in Appendix.
@@
$D$ the discriminator to identify $x$ and $\hat{x}$, a patch-based discriminator Isola et al. (2017)
$\mathcal{L}_{LLM}$ next-token prediction cross-entropy loss with z-loss for training large language model
@@ Appendix C Data Examples
[…]
@@
LogP (Partition Coefficient) measures a molecule’s solubility in fats versus water by quantifying its distribution between octanol (fat-like) and water phases. Calculated as the logarithm of the concentration ratio (LogP = log[octanol]/[water]), it predicts drug absorption and permeability—higher values (>0) indicate greater fat solubility, while lower values (<0) suggest water solubility. Ideal drug candidates typically have LogP between 0-3 for optimal bioavailability. Here is an image of a molecule, please generate an image of a new similar molecule whose LogP is better.[Uncaptioned image] Here is a new similar molecule with better LogP.[Uncaptioned image]
@@ Appendix D Baseline Methods
  • Qwen-VL-Chat Bai et al. (2023) is an open-source multimodal conversational model developed by Alibaba Cloud, extending the Qwen-VL architecture to support complex visual-language interaction through instruction tuning. It integrates a frozen CLIP-ViT-G/14 vision encoder with the Qwen-7B language model via a trainable vision-language connector. Images are encoded into 1024-dimensional patch embeddings by the vision encoder, which are then linearly projected into the token embedding space of the language model (hidden size 4096). The language backbone consists of 32 transformer decoder layers, each employing multi-head masked self-attention with rotary position embeddings (RoPE), followed by a SwiGLU-activated feedforward network with an intermediate dimension of 11008. Qwen-VL-Chat uses a 2048-token context window and supports dynamic multimodal prompts comprising text, images, and region boxes. It is instruction-tuned on a large-scale, GPT-generated multimodal dataset containing both single- and multi-turn visual conversations, enabling capabilities in visual question answering, dense captioning, document OCR, and multi-image reasoning. The model achieves high performance on benchmarks such as MME, SEED-Bench, and MMBench, demonstrating strong alignment between visual and linguistic modalities.
@@
  • InternVL-Chat-V1.5 Chen et al. (2023) is an open-source vision-language instruction-following model developed by OpenGVLab, designed to support high-resolution, multilingual, and multi-turn visual conversations. The model integrates a powerful ViT-based vision encoder, InternViT-6B, with the InternLM2-Chat-20B language model via a trainable multi-layer perceptron (MLP) connector. InternViT encodes images into patch embeddings with dynamic resolution support, allowing the model to process up to 40 image tiles of size 448$\times$448, effectively supporting 4K-level inputs. These embeddings are projected into the language model’s token space to enable seamless multimodal interaction. The language backbone consists of 64 transformer decoder layers with rotary positional embeddings (RoPE), multi-head masked self-attention, and SwiGLU-activated feedforward layers. The model uses a maximum context length of 4096 tokens and is instruction-tuned on a high-quality bilingual dataset containing document images, natural images, and complex multimodal dialogues. InternVL-Chat-V1.5 achieves strong performance on benchmarks including MME, MMBench, and AI2D, demonstrating robust capabilities in visual question answering, document OCR, visual reasoning, and bilingual understanding. The total parameter count is approximately 25.5 billion, with both the vision encoder and language model jointly fine-tuned during the instruction-following phase.
@@
  • LLaVA-v1.5-7B (Large Language and Vision Assistant) Liu et al. (2023) is an open-source vision-language instruction-tuned model that integrates a pre-trained CLIP-ViT-L/14 vision encoder with the LLaMA-7B language model through a projection network. LLaVA-7B processes visual inputs by encoding images into 1024-dimensional patch embeddings via the vision encoder, which are then projected into the language model’s token space through a trainable linear layer (hidden size 4096). The language model comprises 32 transformer decoder blocks, each with masked self-attention (key/query size 4096, 32 attention heads), followed by a feed-forward network with SwiGLU activation and an intermediate dimension of 11008. The self-attention layers use RoPE (rotary positional embeddings) and support a context window of 2048 tokens. LLaVA-7B leverages instruction-tuning on 558K GPT-4 generated multimodal instruction-following samples, aligning visual and textual representations for tasks such as visual QA and image captioning. The total number of trainable parameters is approximately 7 billion, with the vision encoder frozen during fine-tuning.
@@
  • GPT-4o OpenAI (2024) (Generative Pre-trained Transformer 4 Omni) is a state-of-the-art multimodal foundation model developed by OpenAI, designed to natively process and reason across text, images, and audio modalities. Unlike previous GPT-4 variants that rely on separate vision encoders, GPT-4o employs a unified transformer architecture that jointly encodes multimodal inputs, enabling low-latency and high-fidelity interactions. The model supports up to 128k tokens of context and exhibits strong performance across a wide range of tasks, including natural language understanding, image captioning, document analysis, and spoken language comprehension. GPT-4o achieves significant improvements in visual reasoning (e.g., charts, diagrams, OCR), math problem solving, and multilingual capability, surpassing the capabilities of GPT-4-turbo while operating with lower inference latency. It is instruction-tuned on a diverse and extensive corpus of text and multimodal data, and optimized for both conversational fluency and factual grounding. We also utilized GPT-4o to assist in the writing of this paper.
@@
  • ChemLLM-7B-Chat Zhang et al. (2024a) is an open-source domain-specific large language model framework tailored for chemical sciences, designed to address the limitations of general-purpose LLMs in structured scientific domains. The model is instruction-tuned using ChemData, a templated dataset that transforms structured chemical knowledge (e.g., molecular properties, SMILES strings, compound databases) into natural language instructions across a variety of chemical tasks. ChemLLM adopts a decoder-only transformer architecture based on InternLM2-Base-7B Cai et al. (2024), and is fine-tuned to handle both single- and multi-turn dialogues. The model supports tasks such as molecular property prediction, compound generation, synthetic route planning, and reaction condition recommendation.
@@
  • ChemVLM-8B is an open-source multimodal domain-specific large language model specifically designed for chemistry-related tasks, aiming to bridge the gap between vision and language understanding in the chemical domain. The model adopts a ViT-MLP-LLM architecture, integrating a vision encoder based on Vision Transformer (ViT), a multi-layer perceptron (MLP) as a projection layer, and a 20-billion-parameter decoder-only language model ChemLLM-20B Zhang et al. (2024a) as the backbone. ChemVLM processes visual inputs such as molecular structures, chemical reaction schemes, and spectra by encoding images into patch embeddings through ViT, which are then linearly projected to the token space of the language model. It is instruction-tuned on a constructed dataset of 1.2M multimodal samples covering tasks like molecule captioning, reaction classification, and chemical structure understanding.
@@ Appendix E More Visual Result
To better visualize results for different run on five tasks, we draw metric bar for each tasks. The metric bar for img2caption, img2property, img2smiles, property2img, img2img task is shown in Figure 7, Figure 8, Figure 10, Figure 11, Figure 9, respectively.
@@
For the two image generation task, we provide more examples to better display the ability of ChemMLLM. More examples for property2img task are shown in Figure 12. More examples for img2imgtask are shown in Figure 13.

### v2

@@ 6 Conclusion
This paper has proposed ChemMLLM, a chemical multimodal large language model that handles molecule comprehension and generation across text, SMILES string, and molecule image seamlessly. Also, we design five cross-modal chemistry tasks and curate datasets, providing a valuable resource for multimodal AI in chemistry. Across a range of tasks, experimental results demonstrate that ChemMLLMoutperforms state-of-the-art MLLMs and specialized chemical LLMs, highlighting its strong performance and potential for real-world drug and material discovery.
@@ Appendix A Data Curation Details
A.1 Data collection
All the data used in this paper is publicly available and the links are shown in Table 7. In this section, we discuss the details of the data preprocessing steps for each raw datasets.
  • chebi-20 (Edwards et al. 2022): We download training, test and val data in .txt format in the data subfolder, the original data contains 26,407 train data, 3,300 test data and 3,301 val data. Each data entry contains a SMILES-caption pair.
  • Mol-Instructions (Fang et al. 2023): We download description_guided_molecule_design.json and molecular_description_generation.json from Molecule-oriented_Instructions.zip, the original data contains SELFIES-caption pairs with instructions. We first transfer SELFIES into SMILES using RDKit (Landrum et al. 2006), then we filter captions shorter than 150 words to screen clearer descriptions and finally get 43,392 sample.
  • PubChem (Kim et al. 2021): We download all the structure data files (SDF) in the Compound subfolder up to January 13, 2025. Then we use Chem.SDMolSupplier() function in RDKit (Landrum et al. 2006) to parse the original SDF and get 9,691,316 compounds with three fields each related to SMILES, i.e., ”PUBCHEM_SMILES” (the SMILES for compounds), ”PUBCHEM_OPENEYE_CAN_SMILES” (canonical SMILES generated using the OpenEye tool (Software 2023)), and ”PUBCHEM_OPENEYE_ISO_SMILES” (the isomeric smiles generated using the OpenEye tool). After that, we utilize set to remove the duplicate SMILES and finally get 14,874,837 SMILES from raw dataset.
  • TDC (Jin et al. 2018): We download train_pairs.txt and test.txt from logp04 and logp06 subfolder. The original data contains 157,673 train data and 17,520 test data. All data entries contain paired SMILES representations, arranged in ascending order of LogP values (lower LogP $\to$ higher LogP).
@@ A.1 Data collection
Then we utilize the Draw.MolToImage() function in RDKit (Landrum et al. 2006) to transfer SMILES into 256$\times$256 images for all datasets.
Source Link
chebi-20 https://github.com/cnedwards/text2mol
Mol-Instructions https://huggingface.co/datasets/zjunlp/Mol-Instructions
PubChem https://misuse.ncbi.nlm.nih.gov
TDC https://github.com/wengong-jin/iclr19-graph2graph
Table 7: Tasks and datasets.
A.2 Data synthesis
We primarily follow the ” templates filled with raw data” methodology of ChemVLM (Li et al. 2025) to synthesis data.
(1) img2caption: The dataset used for this task is sourced from chebi-20 (Edwards et al. 2022) and Mol-Instructions (Fang et al. 2023). We only use the test set of chebi-20 as the test set and the partitioning of the dataset is the same as chebi-20 (Edwards et al. 2022). The templates for img2caption task can be seen in Table 9.
(2) img2property: The dataset used for this task is sourced from PubChem (Kim et al. 2021). We randomly sample 205,000 SMILES from the original dataset for this work and the random seed is set to 42. And choose 7 important properties as the prediction objective, i.e., MW, LogP, TPSA, HBD, HBA, RB and QED. Then use RDKit to calculate the 7 properties for each sampled SMILES. Then we apply natural language templates to integrate properties into natural language to form the image-property answer pairs. Finally, we divide the dataset into train and test set by the ratio of 40:1. The templates for img2property task can be seen in Table 10.
@@ A.2 Data synthesis
(3) img2smiles: The dataset and the construction steps are the same as the img2property task, the difference is that this task only applies templates for SMILES to construct image-SMILES answer pairs. The templates for img2smiles task can be seen in Table C.
@@
(4) property2img: This task is the inverse of img2property. By swapping the question and answer of the img2property dataset, we construct property prompt-image pairs for property2img task. The templates for property2img task can be seen in Table 12.
(5) img2img: The dataset for this task is sourced from TDC (Jin et al. 2018). For data partitioning, we follow the original dataset. The templates for img2img task can be seen in Table 13.
@@ Appendix B Mathematical Notations
For ease of understanding, we list key mathematical notations in Table 8.
@@
$D$ the discriminator to identify $x$ and $\hat{x}$, a patch-based discriminator (Isola et al. 2017)
$\mathcal{L}$ next-token prediction cross-entropy loss with z-loss for training large language model
@@ Appendix C Data Examples
What is the SMILES of the molecule in this image? [Uncaptioned image]
@@
LogP (Partition Coefficient) measures a molecule’s solubility in fats versus water by quantifying its distribution between octanol (fat-like) and water phases. Calculated as the logarithm of the concentration ratio (LogP = log[octanol]/[water]), it predicts drug absorption and permeability—higher values (¿0) indicate greater fat solubility, while lower values (¡0) suggest water solubility. Ideal drug candidates typically have LogP between 0-3 for optimal bioavailability. Here is an image of a molecule, please generate an image of a new similar molecule whose LogP is better.[Uncaptioned image] Here is a new similar molecule with better LogP.[Uncaptioned image]
@@ Appendix D Baseline Methods
D.1 Task-Specific Models
  • Seq2Seq (Cho et al. 2014) is a foundational sequence-to-sequence architecture designed for statistical machine translation. The model consists of two recurrent neural networks (RNNs): an encoder that reads a variable-length input sequence and compresses it into a fixed-length vector representation, and a decoder that generates a target sequence from this representation.
  • Chemception (Goh et al. 2017) is a convolutional neural network (CNN) model designed for predicting molecular properties using 2D structural images of molecules. Inspired by the Inception architecture, Chemception learns hierarchical features directly from molecular images across various tasks such as toxicity, activity, and solvation energy prediction.
  • Chemprop (Heid et al. 2023) is a graph‑neural‑network based machine learning package designed for molecular property prediction using directed message‑passing neural networks (D‑MPNNs). The model takes molecular graphs (e.g., from SMILES) as input and learns representations by iteratively exchanging messages along directed bonds. Chemprop supports prediction of single‑molecule properties, multimolecule systems, atom‑ and bond‑level targets, and spectral properties.
  • Chemformer (Irwin et al. 2022) is a Transformer‑based model that is pre‑trained in a self‑supervised manner on large corpora of SMILES strings, enabling rapid adaptation to both sequence‑to‑sequence and discriminative cheminformatics tasks like reaction prediction, retrosynthesis planning, molecular optimization, and property classification.
  • MolScribe (Qian et al. 2023) is an image‑to‑graph generation model designed for robust molecular structure recognition from chemical literature figures. The model utilizes a CNN backbone to extract pixel‐level features from the input molecular image, and uses autoregressive Transformer decoder to predict each atom’s label and 2D coordinates, then a fully connected graph classifier infers bond types between atom nodes. The output is assembled into an explicit molecular graph that can be directly used as a node–edge structure or serialized as a SMILES string.
  • DECIMER (Rajan, Zielesny, and Steinbeck 2020) is a deep‑learning based Optical Chemical Structure Recognition (OCSR) model that translates bitmap images of chemical structure depictions into SMILES strings. The model adopts a convolutional backbone (Inception V3) (Szegedy et al. 2016) to encode the input bitmap of a chemical diagram into high‑dimensional feature maps, and utilize a recurrent LSTM decoder to generate the corresponding SMILES string token by token.
D.2 General Multimodal LLM
@@
  • Qwen-VL-Chat (Bai et al. 2023) is an open-source multimodal conversational model developed by Alibaba Cloud, extending the Qwen-VL architecture to support complex visual-language interaction through instruction tuning. It integrates a frozen CLIP-ViT-G/14 vision encoder with the Qwen-7B language model via a trainable vision-language connector. Images are encoded into 1024-dimensional patch embeddings by the vision encoder, which are then linearly projected into the token embedding space of the language model (hidden size 4096). The language backbone consists of 32 transformer decoder layers, each employing multi-head masked self-attention with rotary position embeddings (RoPE), followed by a SwiGLU-activated feedforward network with an intermediate dimension of 11008. Qwen-VL-Chat uses a 2048-token context window and supports dynamic multimodal prompts comprising text, images, and region boxes. It is instruction-tuned on a large-scale, GPT-generated multimodal dataset containing both single- and multi-turn visual conversations, enabling capabilities in visual question answering, dense captioning, document OCR, and multi-image reasoning. The model achieves high performance on benchmarks such as MME, SEED-Bench, and MMBench, demonstrating strong alignment between visual and linguistic modalities.
@@
  • InternVL-Chat-V1.5 (Chen et al. 2023) is an open-source vision-language instruction-following model developed by OpenGVLab, designed to support high-resolution, multilingual, and multi-turn visual conversations. The model integrates a powerful ViT-based vision encoder, InternViT-6B, with the InternLM2-Chat-20B language model via a trainable multi-layer perceptron (MLP) connector. InternViT encodes images into patch embeddings with dynamic resolution support, allowing the model to process up to 40 image tiles of size 448$\times$448, effectively supporting 4K-level inputs. These embeddings are projected into the language model’s token space to enable seamless multimodal interaction. The language backbone consists of 64 transformer decoder layers with rotary positional embeddings (RoPE), multi-head masked self-attention, and SwiGLU-activated feedforward layers. The model uses a maximum context length of 4096 tokens and is instruction-tuned on a high-quality bilingual dataset containing document images, natural images, and complex multimodal dialogues. InternVL-Chat-V1.5 achieves strong performance on benchmarks including MME, MMBench, and AI2D, demonstrating robust capabilities in visual question answering, document OCR, visual reasoning, and bilingual understanding. The total parameter count is approximately 25.5 billion, with both the vision encoder and language model jointly fine-tuned during the instruction-following phase.
@@
  • LLaVA-v1.5-7B (Large Language and Vision Assistant) (Liu et al. 2023) is an open-source vision-language instruction-tuned model that integrates a pre-trained CLIP-ViT-L/14 vision encoder with the LLaMA-7B language model through a projection network. LLaVA-7B processes visual inputs by encoding images into 1024-dimensional patch embeddings via the vision encoder, which are then projected into the language model’s token space through a trainable linear layer (hidden size 4096). The language model comprises 32 transformer decoder blocks, each with masked self-attention (key/query size 4096, 32 attention heads), followed by a feed-forward network with SwiGLU activation and an intermediate dimension of 11008. The self-attention layers use RoPE (rotary positional embeddings) and support a context window of 2048 tokens. LLaVA-7B leverages instruction-tuning on 558K GPT-4 generated multimodal instruction-following samples, aligning visual and textual representations for tasks such as visual QA and image captioning. The total number of trainable parameters is approximately 7 billion, with the vision encoder frozen during fine-tuning.
@@
  • GPT-4o (OpenAI 2024) (Generative Pre-trained Transformer 4 Omni) is a state-of-the-art multimodal foundation model developed by OpenAI, designed to natively process and reason across text, images, and audio modalities. Unlike previous GPT-4 variants that rely on separate vision encoders, GPT-4o employs a unified transformer architecture that jointly encodes multimodal inputs, enabling low-latency and high-fidelity interactions. The model supports up to 128k tokens of context and exhibits strong performance across a wide range of tasks, including natural language understanding, image captioning, document analysis, and spoken language comprehension. GPT-4o achieves significant improvements in visual reasoning (e.g., charts, diagrams, OCR), math problem solving, and multilingual capability, surpassing the capabilities of GPT-4-turbo while operating with lower inference latency. It is instruction-tuned on a diverse and extensive corpus of text and multimodal data, and optimized for both conversational fluency and factual grounding. We also utilized GPT-4o to assist in the writing of this paper.
D.3 Chemical LLM
@@
  • ChemLLM-7B-Chat (Zhang et al. 2024a) is an open-source domain-specific large language model framework tailored for chemical sciences, designed to address the limitations of general-purpose LLMs in structured scientific domains. The model is instruction-tuned using ChemData, a templated dataset that transforms structured chemical knowledge (e.g., molecular properties, SMILES strings, compound databases) into natural language instructions across a variety of chemical tasks. ChemLLM adopts a decoder-only transformer architecture based on InternLM2-Base-7B (Cai et al. 2024), and is fine-tuned to handle both single- and multi-turn dialogues. The model supports tasks such as molecular property prediction, compound generation, synthetic route planning, and reaction condition recommendation.
@@
  • ChemVLM-8B (Li et al. 2025) is an open-source multimodal domain-specific large language model specifically designed for chemistry-related tasks, aiming to bridge the gap between vision and language understanding in the chemical domain. The model adopts a ViT-MLP-LLM architecture, integrating a vision encoder based on Vision Transformer (ViT), a multi-layer perceptron (MLP) as a projection layer, and a 20-billion-parameter decoder-only language model ChemLLM-20B (Zhang et al. 2024a) as the backbone. ChemVLM processes visual inputs such as molecular structures, chemical reaction schemes, and spectra by encoding images into patch embeddings through ViT, which are then linearly projected to the token space of the language model. It is instruction-tuned on a constructed dataset of 1.2M multimodal samples covering tasks like molecule captioning, reaction classification, and chemical structure understanding.
@@ Appendix E More Visual Result
To better visualize results for different run on five tasks, we draw metric bar for each tasks. The metric bar for img2caption, img2property, img2smiles, property2img, img2img task is shown in Figure 9, Figure 10, Figure 12, Figure 13, Figure 11, respectively. The box plots illustrate the performance of different models across several metrics. Each model is represented along the x-axis, while the y-axis shows the corresponding metric values. The plots highlight the distribution of results through their quartiles, with the central line indicating the median and the box representing the interquartile range (IQR).
@@
For the two image generation task, we provide more examples to better display the ability of ChemMLLM. More examples for property2img task are shown in Figure 14. Our model can generate valid molecule images that meet property requirements based on text prompts. More examples for img2img task are shown in Figure 15. Our model can generate valid molecule images with better LogP properties based on the given original molecule images, completing the molecular optimization task in a visual form. Both examples display the ability of our ChemMLLM in generating visuals utilizing LLM in a unified framework.